# Regular black holes and reductions of thermodynamic phase spaces

### v1

@@ III Temperatures of regular black holes
Generally, $T^{0}_{~{}0}=T^{0}_{~{}0}(r,...)$, where ‘…” represents other parameters from the matter fields. After integration, we are left with the relation $M=M(...)$, which imposes an additional constraint between black hole parameters, leading to a reduction of the thermodynamic phase space. It is the chief culprit that leads to the inconsistency of thermodynamic quantities.
@@
which includes only one parameter $\alpha$ that is initially independent of the black hole mass and the magnetic charge. $\alpha$ is a positive coupling constant with the dimension of $[L]^{-2}$ and $F=F_{\mu\nu}F^{\mu\nu}$.
@@ IV Thermodynamics of regular black holes
In Fig.1, it is shown that the thermodynamic behavior of the Bardeen black hole is similar to that of the RN black hole. However, there is still a major difference. The divergent point of the heat capacity does not correspond to the extremal point of the temperature. This correspondence, however, always holds for singular black holes since they satisfy the first law of thermodynamics, where the heat capacity satisfied $C=\frac{\partial M}{\partial T}=\frac{\partial M/\partial r_{+}}{\partial T/\partial r_{+}}$. For the Bardeen black hole, when $T<0.026$, only the smaller black hole has positive heat capacity. When $T>0.026$, the Bardeen black hole admits two locally stable states with identical temperature: a larger black hole phase and a smaller black hole phase. Furthermore, the $F-T$ curve reveals that the smaller black hole phase possesses lower free energy. Consequently, the smaller black hole state is globally thermodynamically stable.
The thermodynamics of regular black holes can also be investigated in asymptotically AdS spacetime, where we expect the possibility of $P-V$ criticality and a richer phase structure.

### v2

@@ III Temperatures of regular black holes
which includes only one parameter $\alpha$ that is initially independent of the black hole mass and the magnetic charge. $\alpha$ is a positive coupling constant with the dimension of $L]^{-2}$ and $F=F_{\mu\nu}F^{\mu\nu}$.
@@ IV Thermodynamics of regular black holes
In Fig.1, it is shown that the thermodynamic behavior of the Bardeen black hole is similar to that of the RN black hole. However, there is still a major difference. The divergent point of the heat capacity does not correspond to the extremal point of the temperature. This correspondence, however, always holds for singular black holes since they satisfy the first law of thermodynamics, where the heat capacity satisfied $C=\frac{\partial M}{\partial T}=\frac{\partial M/\partial r_{+}}{\partial T/\partial r_{+}}$. For the Bardeen black hole, when $T<0.026$, only the smaller black hole has positive heat capacity. When $T>0.026$, the Bardeen black hole admits two locally stable states with identical temperature: a larger black hole phase and a smaller black hole phase. Furthermore, the FTF-T curve reveals that the smaller black hole phase possesses lower free energy. Consequently, the smaller black hole state is globally thermodynamically stable.
The thermodynamics of regular black holes can also be investigated in asymptotically AdS spacetime, where we expect the possibility of PVP-V criticality and a richer phase structure.